Mark the following data on PH chart and calculate Coefficient of performance and heat rejected in...

  1. Mark the following data on PH chart and calculate Coefficient of performance and heat rejected in condenser. The mass flow rate of refrigerant is 0.25 kg/s. The refrigerant used is 134a.          

                              Condenser pressure = 10 bar

                               Evaporator pressure = 0.4 bar.

                              Compressor inlet temperature = -20 ℃

                              Expansion device inlet temperature = 30 ℃                                         
Answer #1

In this problem, first find the specific enthalpy of Refrigerant at each state of the cycle from p-h chart of the refrigerant R-134a and then Calculate the coefficient of performance of cycle and heat rejected in condenser as below:-

Ciren data :-( Refrigerant R-1340) (bar) to- 18= constant (+83k/sk) 0.4 Ft- m P, - PA = 0.4 bar P2 = P3= 10 bar T= -20°C Tz =Copa del Rei CoP = 2:176 Heat rejected in condenser (Qc):- de = mp (he-hs) = 8:25 (456 - 240) kW Te = 54 KW